# Where Quantum Complexity Helps Classical Complexity: A Survey

### v2

@@ 3.4. Quantum Query Complexity
Aaronson and Ambainis (aaronson2015forrelation, ) used a property-testing problem called Forrelation. In this problem, one decides whether a Boolean function is highly correlated with the Fourier transform of a second function. Consider all partial functions [^1] Recall that a partial function is a function not defined for some inputs of the right type. Namely for some of a domain. E.g. division is a partial function since division by 00 is undefined on the Reals. on $N$ bits that can be computed with an advantage $\delta$ over a random guess by making $q$ quantum queries. Aaronson and Ambainis showed that such functions can be computed classically as well, with an advantage $\delta/2$ by a randomized decision tree. In particular, it requires $\mathcal{O}_{q}(N^{1-\frac{1}{2q}}\delta^{-2})$ many queries. Let the k-Forrelation problem be a partial function taking $q=\lceil k/2\rceil$ quantum queries. The authors in (aaronson2015forrelation, ) conjectured that k-Forrelation is a suitable candidate for exhibiting such an extremal separation.
Bansal and Sinha (bansal2021k, ) consider the most basic access model in their work. In particular, each query returns a bit of the input. They prove the conjecture of Aaronson and Ambainis by showing a tight lower bound of $\tilde{\Omega}(N^{1-\frac{1}{k}})$ for the randomized query complexity of k-Forrelation s.t. $\delta=2^{-\mathcal{O}(k)}$.
Consider a classical query algorithm (decision tree) and a guessing algorithm that tries to predict the query answers. One may then suggest a quantum algorithm with query complexity $\mathcal{O}(\sqrt{GT})$ s.t. $T$ denotes the query complexity of the classical algorithm (i.e. depth of the decision tree). Also, $G$ represents the maximum number of guessing algorithm’s wrong answers (Beigi2020quantumspeedupbased, ; v012a018, ). Beigi et al. (beigi2022time, ) proved that, provided some constraints on the classical algorithms, one may implement this quantum algorithm in time $\tilde{\mathcal{O}}(\sqrt{GT})$. The algorithm in (beigi2022time, ) is based on non-binary span programs and their efficient implementation. Let $n$ be the number of vertices of a given graph. The authors pointed out that various graph-theoretic problems, including bipartiteness, cycle detection, and topological sort can be solved in time $\mathcal{O}(n^{3/2}\log^{2}n)$ and with $\mathcal{O}(n^{3/2})$ quantum queries. Further, one may find a maximal matching with $\mathcal{O}(n^{3/2})$ quantum queries in time $\mathcal{O}(n^{3/2}\log^{2}n)$, and maximum bipartite matching in time $\mathcal{O}(n^{2}\log^{2}n)$.
@@ 3.5. Quantum Random Walk
Apers et al. (apers2019unified, ) introduced a quantum walk search framework, unifying and strengthening quantum walk frameworks. Their framework detects and finds marked elements in the electric network setting. It also permits interpolating between the hitting time framework. This minimizes the number of walk steps, and the MNRS framework (magniez2007search, ), minimizes the number of checking the elements for being marked. The previous frameworks relied on quantum walks and phase estimation only. Nonetheless, the algorithm used in (apers2019unified, ) benefits from a technique named quantum fast-forwarding as well. Further, the authors showed how in certain cases one may perform simplifications by applying the quantum walk operator several times.
The shortest vector problem (SVP) is an important problem for the cryptanalysis of lattice-based cryptography. Chailloux and Loyer (chailloux2021lattice, ) improved the results of Laarhoven (laarhoven2016search, ) for SVP. They introduced an algorithm with a (heuristic) running time of $2^{0.2570d+o(d)}$ s.t. $d$ is the lattice dimension. The authors also suggested time-memory trade-offs. Namely, one may quantify the amount of quantum memory and quantum random access memory of the algorithm. From a broader perspective, the idea of their work is to replace Grover’s algorithm used in (laarhoven2016search, ) in a key part of the sieving algorithm. This is done by a quantum random walk in which one adds a layer of local sensitive filtering.
Marsh and Wang (marsh2019quantum, ) described an application of QAOA. They aimed to find approximate solutions for computational problems contained in NPO PB222polynomially bounded NP optimization complexity class efficiently. The algorithm QAOA is then applied by the authors to the minimum vertex cover problem. They studied a generalization of the QAOA state evolution to alternating quantum walks and solution-quality-dependent phase shifts. They also use quantum walks to integrate the problem limitations of NPO problems. Marsh and Wang (marsh2019quantum, ) have shown that a hybrid quantum-classical variational algorithm suits finding the value of a certain expectation value. It is also efficient for NP optimization problems that have polynomially bounded measures. The results in (marsh2019quantum, ) illustrated that for solving the minimum vertex problem, a hybrid quantum-classical variational scheme shows promising results. It only requires hiring a fixed and small number of optimization parameters.
@@
Consider the welded trees problem. The authors in (jeffery2022multidimensional, ) suggested a solution requiring $\mathcal{O}(n)$ queries and $\mathcal{O}(n^{2})$ time. The result is a direct application of the technique mentioned earlier. Namely, the quantum walk framework suggested in (jeffery2022multidimensional, ) achieves exponential improvement in terms of running time.
@@ 4. Quantum Solutions to Classical Problems
Quantum annealers (QAs) physically exploit quantum effects. They are specialized quantum computers that minimize objective functions over discrete variables. Many QA platforms allow the optimization of quadratic objectives defined over binary variables (i.e. qubits) a.k.a Ising Problems. The QA systems implemented by D-Wave followed a Moore-like growth in terms of scaling. There are architectures with $2048$ sparsely-connected qubits. Continued exponential growth and increased connectivity are anticipated.
@@ 4.1. Satisfiability Problem
Classical algorithms were previously applied to solve the satisfiability problem. Due to the insufficient computational power of classical computers, effective methods are still being explored. Quantum algorithms are among the options, as they have the ability of parallelism. Nonetheless, many quantum algorithms require a large number of qubits to solve a computationally simpler problem. Zhang et al. (zhang2020procedure, ) suggested an optimized data structure to solve the Boolean satisfiability problem. The authors utilized Grover’s algorithm first. Second, they proposed a formula, balancing variables in consideration of complexity. Via simplification, the authors built quantum circuits to decrease the number of qubits required in Grover’s algorithm.
@@
Cheng et al. (CHENG2007123, ) propose a new quantum cooperative search algorithm to make Grover’s search algorithm work with a small number of quantum bits in solving the 3-SAT problem. The proposed algorithm, in use of traditional 3-SAT algorithms such as evolutionary algorithms and heuristics local search algorithms, is replacing some quantum bits with classical bits to identify assignments for these classical bits. Moreover, a mathematical analysis suggests the optimal configuration of the proposed algorithm. In most cases, the experimental results show that a quantum cooperative search algorithm composed of Grover’s search and heuristic local search performs better than other pure traditional 3SAT algorithms. Furthermore, the experiments confirm a mathematical analysis of an appropriate number of quantum bits (CHENG2007123, ).
@@
Varmantchaonala et al. (varmantchaonala2023quantum, ) studied a particular parameter. The authors suggested an approach based on the Grover algorithm. Their approach considers the aforementioned parameter, reducing the search space for the SAT problem. Let $l$ be the number of good solutions amongst the initial solutions. Then the algorithm in (varmantchaonala2023quantum, ) requires the time complexity of $\mathcal{O}(\sqrt{N/lS})$ s.t. $N$ is the number of potential solutions. Also, $S$ denotes the reduction or subdivision factor of the space. In case there are no such $l$ solutions, the time complexity becomes proportional to $\mathcal{O}(\sqrt{N/S})$.
Leporati and Felloni (leporati2007three, ) proposed three quantum algorithms that solve the 3-SAT decision problem in polynomial time, in the semi-uniform setting. Their construction assumes an external observer may measure a null vector and a non-null vector.
Consider every instance $\phi_{n}$ of 3-SAT. The first algorithm in (leporati2007three, ) constructs a quantum Fredkin circuit. This circuit computes a superposition of all classical evaluations of $\phi_{n}$ in the first output line.
@@
Given a register of a quantum register machine. Similar to the first one, the second and third algorithms compute the same superposition on the register, and as the energy of a given membrane in a quantum $P$ system, respectively. Provided a non-unitary operator. Assume that the operator can be realized as a quantum gate. That is, an instruction of the quantum register machine, and as a rule of the quantum $P$ system. The operator can be described using creation and annihilation operators. As stated in (leporati2007three, ), one may apply the operator to the result of the computation performed by the algorithms. So, the solution of 3-SAT for the instance $\phi_{n}$ can be extracted.
Conflict-Driven Clause Learning (CDCL) is an algorithm for solving SAT. In the case of a Boolean formula, the SAT problem requires the assignment of variables so that the entire formula can be evaluated as true.
Quantum annealing (QA) is a process of optimization for finding the global minimum of a function over a set of solutions using quantum fluctuations.
@@
Tan et al. (10071022, ) proposed HyQSAT, a hybrid approach that integrates QA with Conflict-Driven Clause Learning. It enables end-to-end acceleration for solving SAT problems. Instead of all embedded clauses for QA hardware, the authors used a quantitative estimation of conflict frequency and breadth-first traversal to determine their embedded order. HyQSAT takes advantage of both quantum parallelism and classical algorithms. Further, the experiments in (10071022, ) showed that, compared to the state of classical physics, HyQSAT can achieve a real quantum speed of 12.62X on DWave 2000s.
Ohya-Masuda Quantum algorithm uses a quantum computer and a classical amplifier for solving the 3-SAT problem (ohya1999quantum, ). The 3-SAT may be solved in polynomial time by quantum computer if the superposition of two orthogonal vectors |$\ket{0}$ and |$\ket{1}$ can physically be detected (ohya1998np, ). Nonetheless, the issue is that it is inconvenient to distinguish the items with a low measurement probability.
@@
The algorithm in (ohya1999quantum, ) consists of three steps. The first one performs a unitary transformation on the input state. The second one carries out a nonunitary transformation that gives a classical mixture of states. Lastly, the third step represents a logistic map amplification of the result of the second step. However, Dugic in (10.1023/A:1015644508998, ) described that the crucial second step is not possible to be achieved in general. The heart of the Ohya-Masuda algorithm is producing a classical mixture out of pure states. Then, amplifying classical information on the last qubit state. In the second step of the algorithm, a non-unitary transformation is used. In a realistic situation, it is unlikely that such a transformation will occur. And, in most cases, an error related to an imperfect measurement will change the final state unavoidably. So, this error and its amplification give the wrong result in the end (10.1023/A:1015644508998, ).
@@
Benjamin et al. (benjamin2017measurement, ) investigated how a quantum algorithm performs when solving the 3-SAT problem. The authors point out that a cycle of post-selected measurements drives monotonically the computer’s register toward a steady state. This implies a correlation to the classical solution(s). Let $\theta$ be an internal parameter that determines the degree of correlation and the probability of success. So, $\theta$ controls the algorithm’s runtime. $\theta$ can optionally be evolved gradually during the algorithm’s execution. Evolving the parameter then creates a Zeno-like effect. One may view this as an adiabatic evolution of a Hamiltonian s.t. it remains frustration-free. It lower bounds the corresponding gap as well. Consider using up to thirty-four qubits when performing the exact numerical simulations of small systems. The approach in (benjamin2017measurement, ) offers a competitive running time compared to a classical 3-SAT solver. A classical 3-SAT solver itself outperforms a brute-force approach i.e. applying Grover’s search.
@@
Authors in (nusslein2023solving, ) introduced an approach to translate arbitrary 3-SAT instances to QUBO. Such 3-SAT instances are used by quantum annealing or the quantum approximate optimization algorithm. The approach in (nusslein2023solving, ) requires a smaller number of couplings and fewer physical qubits than its previous methods. This advantage results in higher solution quality. In addition, the authors verified the benefits of their approach in practice by testing it on a D-Wave quantum annealer.
@@ 4.2. Maximum Independent Set
An internal set problem is a graph optimization problem that involves determining subsets of vertices on the graph, in which there are no two vertices connected to one another.
@@
In the above definitions, $V$ is the set of vertices of which a component $(j,k)$ corresponds to the literal $l_{j,k}$ in the $j$-th clause $C_{j}$. Also, $E$ is the set of all edges, namely the union of two edge sets $E_{1}$ and $E_{2}$. In particular, $E_{1}$ connects vertices in the same clause and $E_{2}$ connects vertices in different clauses negating each other. Figure 1 is an instance of the MIS graph.
@@
In the above equation, $U$ is the van der Walls interaction between nearest-neighbor in fixed distance $d$, $\delta$ is the time-dependent laser detuning, and $\Omega$ is the time-dependent Rabi frequency (kim2022rydberg, ). For short, laser detuning is adjusting the laser to a slightly different frequency than the frequency of the quantum system. Also, Rabi frequency is the frequency at which the energy level of an atom fluctuates in an electromagnetic field.
Dörn (doern2007quantum, ) presented a reduced and upper limit of quantum complexity in terms of independent set problems for graphs. Specifically, the author provided quantum algorithms to compute a graph’s maximal and maximum independent sets. The results have improved the best classical bounds of complexity for some graph problems (doern2007quantum, )
@@
Consider the UnitDisk Maximum Independent Set problem. Serret et al. (Serret_2020, ) calculated the quantified requirements on system size and noise levels that platforms must fulfill to achieve quantum benefits in solving the problem. The authors computed the average approximation ratio, using noisy simulations of Rydberg platforms of up to twenty-six atoms. The atoms interacted through realistic van der Waals interactions. The above can be attained with a simple quantum annealing-based heuristic within a fixed temporal computational budget. For near-future noise levels, the authors found that ratios of at least $0.84$ can be reached for approximation. Consequently, quantum advantages can be obtained by two possible hardware-related conditions: $(i)$ An increase in the size of the quantum system to a few thousand atoms; $(ii)$ An improvement in the quantum system’s coherence properties. The authors emphasized that one might achieve a higher quantitative approximation ratio by pure algorithmic improvements, namely, beyond the simple quantum annealing method used in (Serret_2020, ).
@@
Finding an independent set with the largest possible size for a given graph is required to solve the maximum independent set problem. It is known that this problem is NP-hard, which means that there is unlikely to be an efficient algorithm to find the maximum independent set of graphs. The authors provided the first results of experiments involving arbitrarily structured MIS inputs using a D-Wave QPU (Commercial quantum processing unit), which is increasingly used to solve complex combinatorial optimization problems. The approach of Yarkoni et al. (8477865, ) shows how one may map MIS to QUBO problems. In their work, the authors had a limiting factor, the number of QPU (Quantum Processing Unit) parameters to be tuned to optimize performance (8477865, ).

### v4

@@ 4.4. Quantum Query Complexity
Aaronson and Ambainis (aaronson2015forrelation, ) used a property-testing problem called Forrelation. In this problem, one decides whether a Boolean function is highly correlated with the Fourier transform of a second function. Consider all partial functions [^1] Recall that a partial function is a function not defined for some inputs of the right type. Namely for some of the domains. E.g. division is a partial function since division by 00 is undefined on the Reals. on $N$ bits that can be computed with an advantage $\delta$ over a random guess by making $q$ quantum queries. Aaronson and Ambainis showed that such functions can be computed classically as well, with an advantage $\delta/2$ by a randomized decision tree. In particular, it requires $\mathcal{O}_{q}(N^{1-\frac{1}{2q}}\delta^{-2})$ many queries. Let the k-Forrelation problem be a partial function taking $q=\lceil k/2\rceil$ quantum queries. The authors in (aaronson2015forrelation, ) conjectured that k-Forrelation is a suitable candidate for exhibiting such an extremal separation.
Further, Bansal and Sinha (bansal2021k, ) considered the most basic access model in their work. In particular, each query returns a bit of the input. They proved the conjecture of Aaronson and Ambainis by showing a tight lower bound of $\tilde{\Omega}(N^{1-\frac{1}{k}})$ for the randomized query complexity of k-Forrelation s.t. $\delta=2^{-\mathcal{O}(k)}$.
Apart from the works related to the Forrelation, consider a classical query algorithm (decision tree) and a guessing algorithm that tries to predict the query answers. One may then suggest a quantum algorithm with query complexity $\mathcal{O}(\sqrt{GT})$ s.t. $T$ denotes the query complexity of the classical algorithm (i.e. depth of the decision tree). Also, $G$ represents the maximum number of guessing algorithm’s wrong answers (Beigi2020quantumspeedupbased, ; v012a018, ). Beigi et al. (beigi2022time, ) proved that, provided some constraints on the classical algorithms, one may implement this quantum algorithm in time $\tilde{\mathcal{O}}(\sqrt{GT})$. The algorithm in (beigi2022time, ) is based on non-binary span programs and their efficient implementation. Let $n$ be the number of vertices of a given graph. The authors pointed out that various graph-theoretic problems, including bipartiteness, cycle detection, and topological sort can be solved in time $\mathcal{O}(n^{3/2}\log^{2}n)$ and with $\mathcal{O}(n^{3/2})$ quantum queries. Further, one may find a maximal matching with $\mathcal{O}(n^{3/2})$ quantum queries in time $\mathcal{O}(n^{3/2}\log^{2}n)$, and maximum bipartite matching in time $\mathcal{O}(n^{2}\log^{2}n)$.
@@ 4.5. Quantum Random Walk
Apers et al. (apers2019unified, ) introduced a quantum walk search framework, unifying and strengthening quantum walk frameworks. Their framework detects and finds marked elements in the electric network setting. It also permits interpolating between the hitting time framework. This minimizes the number of walk steps, and the MNRS framework (magniez2007search, ), minimizes the number of checking the elements for being marked. The previous frameworks relied on quantum walks and phase estimation only. Nonetheless, the algorithm used in (apers2019unified, ) benefits from a technique named quantum fast-forwarding as well. Further, the authors showed how, in certain cases, one may perform simplifications by applying the quantum walk operator several times.
Another interesting application of quantum random walk is in the shortest vector problem (SVP). Which, is an important problem for the cryptanalysis of lattice-based cryptography. Chailloux and Loyer (chailloux2021lattice, ) improved the results of Laarhoven (laarhoven2016search, ) for SVP. They introduced an algorithm with a (heuristic) running time of $2^{0.2570d+o(d)}$ s.t. $d$ is the lattice dimension. The authors also suggested time-memory trade-offs. Namely, one may quantify the amount of quantum memory and quantum random access memory of the algorithm. From a broader perspective, their idea is to replace Grover’s algorithm used in (laarhoven2016search, ) in a key part of the sieving algorithm. This is done by a quantum random walk in which one adds a layer of local sensitive filtering.
Marsh and Wang (marsh2019quantum, ) aimed to find approximate solutions for computational problems contained in NPO PB222polynomially bounded NP optimization complexity class efficiently. The authors then apply QAOA to the minimum vertex cover problem. They studied a generalization of the QAOA state evolution to alternating quantum walks and solution-quality-dependent phase shifts. They also use quantum walks to integrate the problem limitations of NPO problems. Marsh and Wang (marsh2019quantum, ) have shown that a hybrid quantum-classical variational algorithm suits finding the value of a certain expectation value. It is also efficient for NP optimization problems that have polynomially bounded measures. The results in (marsh2019quantum, ) illustrated that for solving the minimum vertex problem, a hybrid quantum-classical variational scheme shows promising results. It only requires hiring a fixed and small number of optimization parameters.
@@
Consider the welded trees problem. The authors in (jeffery2022multidimensional, ) suggested a solution requiring $\mathcal{O}(n)$ queries and $\mathcal{O}(n^{2})$ time. The result is a direct application of the technique mentioned earlier. Namely, the quantum walk framework suggested in (jeffery2022multidimensional, ) achieves exponential improvement in running time.
@@ 5. Quantum Solutions to Classical Problems
Further, Quantum Annealers (QAs) physically exploit quantum effects. They are specialized quantum computers that minimize objective functions over discrete variables. Many QA platforms allow the optimization of quadratic objectives defined over binary variables (i.e. qubits) a.k.a Ising Problems. The QA systems implemented by D-Wave followed a Moore-like growth in terms of scaling. There are architectures with $2048$ sparsely-connected qubits. Continued exponential growth and increased connectivity are anticipated.
@@ 5.1. Satisfiability Problem
Classical algorithms were previously applied to solve the satisfiability problem. Due to the insufficient computational power of classical computers, effective methods are still being explored. Quantum algorithms are among the options, as they can parallelism. Nonetheless, many quantum algorithms require a large number of qubits to solve a computationally simpler problem.
Zhang et al. (zhang2020procedure, ) suggested an optimized data structure to solve the Boolean satisfiability problem. They utilized Grover’s algorithm first. Second, they proposed a formula, balancing variables in consideration of complexity. Via simplification, the authors built quantum circuits to decrease the number of qubits required in Grover’s algorithm.
@@
Cheng et al. (CHENG2007123, ) proposed a new quantum cooperative search algorithm to make Grover’s search algorithm work with a small number of quantum bits in solving the 3-SAT problem. The proposed algorithm, in the use of traditional 3-SAT algorithms such as evolutionary algorithms and heuristics local search algorithms, replaces some quantum bits with classical bits to identify assignments for these classical bits. Moreover, a mathematical analysis suggests the optimal configuration of the proposed algorithm. In most cases, the experimental results show that a quantum cooperative search algorithm composed of Grover’s search and heuristic local search performs better than other pure traditional 3SAT algorithms. Furthermore, the experiments confirm a mathematical analysis of an appropriate number of quantum bits (CHENG2007123, ).
@@
Varmantchaonala et al. (varmantchaonala2023quantum, ) studied a particular parameter. They suggested an approach based on the Grover algorithm. Their approach considers the aforementioned parameter, reducing the search space for the SAT problem. Let $l$ be the number of good solutions amongst the initial solutions. Then the algorithm in (varmantchaonala2023quantum, ) requires the time complexity of $\mathcal{O}(\sqrt{N/lS})$ s.t. $N$ is the number of potential solutions. Also, $S$ denotes the reduction or subdivision factor of the space. In case there are no such $l$ solutions, the time complexity becomes proportional to $\mathcal{O}(\sqrt{N/S})$.
@@
Leporati and Felloni (leporati2007three, ) proposed three quantum algorithms that solve the 3-SAT decision problem in polynomial time, in the semi-uniform setting. Their construction assumes an external observer may measure a null vector and a non-null vector. Consider every instance $\phi_{n}$ of 3-SAT. The first algorithm in (leporati2007three, ) constructs a quantum Fredkin circuit. This circuit computes a superposition of all classical evaluations of $\phi_{n}$ in the first output line. Now, suppose a register of a quantum register machine is given. Similar to the first algorithm, the second and third algorithms compute the same superposition on the register, and as the energy of a given membrane in a quantum $P$ system, respectively. Provided a non-unitary operator. Assume that the operator can be realized as a quantum gate. That is, an instruction of the quantum register machine, and as a rule of the quantum $P$ system. The operator can be described using creation and annihilation operators. As stated in (leporati2007three, ), one may apply the operator to the result of the computation performed by the algorithms. So, the solution of 3-SAT for the instance $\phi_{n}$ can be extracted.
@@
Moving to another method, consider Conflict-Driven Clause Learning (CDCL), that is, an algorithm for solving SAT. In the case of a Boolean formula, the SAT problem requires the assignment of variables so that the entire formula can be evaluated as true. Tan et al. (10071022, ) proposed HyQSAT, a hybrid approach that integrates QA with Conflict-Driven Clause Learning. It enables end-to-end acceleration for solving SAT problems. Instead of all embedded clauses for QA hardware, the authors used a quantitative estimation of conflict frequency and breadth-first traversal to determine their embedded order. HyQSAT takes advantage of both quantum parallelism and classical algorithms. Further, the experiments in (10071022, ) showed that, compared to the state of classical physics, HyQSAT can achieve a real quantum speed of 12.62X on DWave 2000s.
@@
Yet, another method is called the Ohya-Masuda Quantum algorithm, which uses a quantum computer and a classical amplifier for solving the 3-SAT problem (ohya1999quantum, ). The 3-SAT may be solved in polynomial time by quantum computer if the superposition of two orthogonal vectors |$\ket{0}$ and |$\ket{1}$ can physically be detected (ohya1998np, ). Nonetheless, the issue is that it is inconvenient to distinguish the items with a low measurement probability. The algorithm in (ohya1999quantum, ) consists of three steps. The first one performs a unitary transformation on the input state. The second one carries out a nonunitary transformation that gives a classical mixture of states. Lastly, the third step represents a logistic map amplification of the result of the second step. However, Dugic in (10.1023/A:1015644508998, ) described that the crucial second step is not possible to be achieved in general. The heart of the Ohya-Masuda algorithm is producing a classical mixture out of pure states. Then, amplifying classical information on the last qubit state. In the second step of the algorithm, a non-unitary transformation is used. In a realistic situation, it is unlikely that such a transformation will occur. And, in most cases, an error related to an imperfect measurement will change the final state unavoidably. So, this error and its amplification give the wrong result in the end (10.1023/A:1015644508998, ).
@@
Apart from the above methods, Benjamin et al. (benjamin2017measurement, ) investigated how a quantum algorithm performs when solving the 3-SAT problem. The authors point out that a cycle of post-selected measurements drives monotonically the computer’s register toward a steady state. This implies a correlation to the classical solution(s). Let $\theta$ be an internal parameter that determines the degree of correlation and the probability of success. So, $\theta$ controls the algorithm’s runtime. $\theta$ can optionally be evolved gradually during the algorithm’s execution. Evolving the parameter then creates a Zeno-like effect. One may view this as an adiabatic evolution of a Hamiltonian s.t. it remains frustration-free. It lower bounds the corresponding gap as well. Consider using up to thirty-four qubits when performing the exact numerical simulations of small systems. The approach in (benjamin2017measurement, ) offers a competitive running time compared to a classical 3-SAT solver. A classical 3-SAT solver itself outperforms a brute-force approach i.e. applying Grover’s search.
@@
Another interesting result is obtained by Nüßlein et al. (nusslein2023solving, ), where they introduced an approach to translate arbitrary 3-SAT instances to QUBO. Such 3-SAT instances are used by quantum annealing or the quantum approximate optimization algorithm. The approach in (nusslein2023solving, ) requires a smaller number of couplings and fewer physical qubits than its previous methods. This advantage results in higher solution quality. In addition, the authors verified the benefits of their approach in practice by testing it on a D-Wave quantum annealer.
@@ 5.2. Maximum Independent Set
An internal set problem is a graph optimization problem that involves determining subsets of vertices on the graph, in which there are no two vertices connected.
@@
such that $V$ is the set of vertices of which a component $(j,k)$ corresponds to the literal $l_{j,k}$ in the $j$-th clause $C_{j}$. Also, $E$ is the set of all edges, namely the union of two edge sets $E_{1}$ and $E_{2}$. In particular, $E_{1}$ connects vertices in the same clause and $E_{2}$ connects vertices in different clauses negating each other. Figure 5 is an instance of the MIS graph.
@@
In the above equation, $U$ is the van der Walls interaction between nearest-neighbour in fixed distance $d$, $\delta$ is the time-dependent laser detuning, and $\Omega$ is the time-dependent Rabi frequency (kim2022rydberg, ). For short, laser detuning is adjusting the laser to a slightly different frequency than the frequency of the quantum system. Also, Rabi frequency is the frequency at which the energy level of an atom fluctuates in an electromagnetic field.
Dörn (doern2007quantum, ) presented a reduced and upper limit of quantum complexity in terms of independent set problems for graphs. Specifically, the author provided quantum algorithms to compute a graph’s maximal and maximum independent sets. The results have improved the best classical bounds of complexity for some graph problems (doern2007quantum, ).
@@
Serret et al. (Serret_2020, ) calculated the quantified requirements on system size and noise levels that platforms must fulfil to achieve quantum benefits in solving the problem. The authors computed the average approximation ratio, using noisy simulations of Rydberg platforms of up to twenty-six atoms. The atoms interacted through realistic van der Waals interactions. The above can be attained with a simple quantum annealing-based heuristic within a fixed temporal computational budget. For near-future noise levels, the authors found that ratios of at least $0.84$ can be reached for approximation. Consequently, quantum advantages can be obtained by two possible hardware-related conditions: $(i)$ An increase in the size of the quantum system to a few thousand atoms; $(ii)$ An improvement in the quantum system’s coherence properties. The authors emphasized that one might achieve a higher quantitative approximation ratio by pure algorithmic improvements, namely, beyond the simple quantum annealing method used in (Serret_2020, ).
@@
Finding an independent set with the largest possible size for a given graph is required to solve the maximum independent set problem. It is known that this problem is NP-hard, which means that an efficient algorithm is unlikely to find the maximum independent set of graphs. The authors provided the first results of experiments involving arbitrarily structured MIS inputs using a D-Wave QPU (Commercial quantum processing unit), which is increasingly used to solve complex combinatorial optimization problems. The approach of Yarkoni et al. (8477865, ) shows how one may map MIS to QUBO problems. In their work, the authors had a limiting factor, the number of QPU (Quantum Processing Unit) parameters to be tuned to optimize performance (8477865, ).